on a map the scale is 1/2 inches =25 miles and I k=need to find the actual distance. Can you help me find the actual distance for 8 and 3/4inches?

Found 2 solutions by checkley79, jim_thompson5910:
Answer by checkley79(3341)

.5/8.75=25/X CROSS MULTIPLY.
.5X=8.75*25
.5X=218.75
X=218.75/.5
X=437.5 MILES.

Answer by jim_thompson5910(35256)

8 and 3/4 = 8 + 3/4

8 and 3/4 = 32/4 + 3/4

8 and 3/4 = 35/4
------------------------------------------

1/2 inches = 25 miles

1 inch = 2*25 miles

1 inch = 50 miles

35/4 inches = (35/4)*50 miles

35/4 inches = (35*50)/4 miles

35/4 inches = 1750/4 miles

35/4 inches = 875/2 miles

35/4 inches = (874 + 1)/2 miles

35/4 inches = 874/2 + 1/2 miles

35/4 inches = 437 + 1/2 miles

35/4 inches = 437 and 1/2 miles

8 and 3/4 inches = 437 and 1/2 miles